# KBQA-R1: Reinforcing Large Language Models for Knowledge Base Question Answering

### v1

@@ Abstract
Knowledge Base Question Answering (KBQA) challenges models to bridge the gap between natural language and strict knowledge graph schemas by generating executable logical forms. While Large Language Models (LLMs) have advanced this field, current approaches often struggle with a dichotomy of failure: they either generate hallucinated queries without verifying schema existence or exhibit rigid, template-based reasoning that mimics synthesized traces without true comprehension of the environment. To address these limitations, we present KBQA-R1, a framework that shifts the paradigm from text imitation to interaction optimization via Reinforcement Learning. Treating KBQA as a multi-turn decision process, our model learns to navigate the knowledge base using a list of actions, leveraging Group Relative Policy Optimization (GRPO) to refine its strategies based on concrete execution feedback rather than static supervision. Furthermore, we introduce Referenced Rejection Sampling (RRS), a data synthesis method that resolves cold-start challenges by strictly aligning reasoning traces with ground-truth action sequences. Extensive experiments on WebQSP, GrailQA, and GraphQuestions demonstrate that KBQA-R1 achieves state-of-the-art performance, effectively grounding LLM reasoning in verifiable execution.
@@ I Introduction
Despite significant progress in applying LLMs to KBQA, existing methodologies can be categorized into three paradigms, each with distinct limitations. The first category comprises End-to-end Approaches (e.g., KB-BINDER [KB-BINDER], KB-Coder [KB-Coder], ChatKBQA [ChatKBQA]), which generate entire logical forms in a single pass. While computationally efficient, these methods operate without intermediate KB interaction, making them unable to verify the existence of schema elements during generation. This often leads to schema hallucinations—syntactically valid but semantically incorrect queries that reference non-existent relations or entities. The second category, Prompting-based Step-by-Step Approaches [ToG, RoG], leverages powerful commercial model APIs with few-shot in-context learning to decompose complex queries into sequential reasoning steps. While these methods benefit from the strong reasoning capabilities of large-scale models, they lack task-specific training, resulting in suboptimal performance on domain-specific schema navigation and complex multi-hop queries. The third category, Supervised Agent Approaches [luo2025kbqa], mitigates the above issues by fine-tuning models on reasoning traces synthesized from templates or heuristics. However, this paradigm risks superficial reasoning: since the training data is inherently formulaic, the model’s “thoughts” often reduce to template-driven action announcements (e.g., “At this step, we should find the relation…”) rather than genuine analysis of environmental feedback. The model declares what action to take without explaining why—it neither interprets the KB observations nor justifies its choices based on the question semantics. Such rigid patterns, while achieving local consistency, fail to generalize when novel schema structures or unexpected query patterns arise.
To address these limitations, we present KBQA-R1, an action-centric reinforcement learning framework that shifts the paradigm from text imitation to interaction optimization. Instead of generating raw query code, KBQA-R1 operates within a well-defined action space, treating KBQA as a multi-turn Markov Decision Process (MDP). At each step, the model selects an action (e.g., Find_Relation, Merge), observes concrete feedback from the KB execution engine, and dynamically adjusts its reasoning trajectory. Crucially, because the policy is optimized via Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] with outcome-based rewards rather than imitation loss, the model is incentivized to develop adaptive reasoning—genuinely analyzing observations and justifying action choices—rather than memorizing fixed reasoning templates. This closed-loop interaction grounds decisions in verifiable outcomes and enables the model to discover effective reasoning strategies through environmental exploration.
@@ II Related Work
Knowledge Base Question Answering (KBQA). Before the rise of LLMs, KBQA studies are commonly categorized into information-retrieval-based (IR-based) methods [GRAFT-Net, PullNet, SR, NSM, EmbedKGQA] and semantic-parsing-based (SP-based) methods [RnG-KBQA, TIARA, ArcaneQA, FC-KBQA]. With LLMs, three paradigms have emerged: (i) end-to-end approaches that directly generate logical forms via in-context learning or fine-tuning [KB-BINDER, KB-Coder, ChatKBQA, StructGPT]; (ii) step-by-step (agentic) approaches that interleave reasoning with graph exploration and tool use [Pangu, QueryAgent, ToG, RoG, Interactive-KBQA, PoG, KG-Agent]; and (iii) search-augmented approaches that leverage tree search algorithms such as Monte Carlo Tree Search (MCTS) for systematic exploration [luo2025kbqa].
@@
While MCTS-based methods like KBQA-o1 [luo2025kbqa] achieve strong performance through heuristic exploration, they exhibit two key limitations. First, they incur significant computational overhead from multiple rollouts per query and require separate policy and reward models during inference. Second, their reasoning traces are often template-driven (e.g., “At this step, we should find the relation…”) rather than genuinely analytical—the model announces what action to take without explaining why based on observations. In contrast, we train a single policy via RL with outcome-based rewards, encouraging the model to develop adaptive reasoning that analyzes environmental feedback and justifies action choices, while eliminating test-time search overhead.
@@
LLMs, tool use, and agentic reasoning. Chain-of-Thought (CoT) prompting improves reasoning by eliciting intermediate steps [CoT]; ReAct [ReAct] interleaves “think” and “act” to ground reasoning in environment feedback; and heuristic search has been applied to agent traces (e.g., MCTS-style selection in [RAP] and tree-structured deliberation in [ToT]). Recent graph-augmented approaches such as Plan-on-Graph [PoG] incorporate self-correcting mechanisms with dynamic memory for adaptive planning on knowledge graphs. While these methods expand the search space or stabilize multi-step reasoning, free-form thoughts can overfit prompt templates and do not guarantee executability. We keep the interleaved think-act design but require typed, schema-aware actions with validators and an executor, turning traces into verifiable computations rather than narrative justifications.
@@
Retrieval-augmented generation and search-as-a-tool. Classical RAG pipelines retrieve text snippets and feed them to the model for generation [RAG]. Recent work moves toward search-as-a-tool, prompting or training LLMs to issue search calls and iterate [trivedi2022interleaving, ReAct, schick2023toolformer, li2025search, jin2025search]. GraphRAG approaches [G-Retriever, ToG2] further integrate graph retrieval with LLM reasoning, enabling tighter coupling between structured knowledge and text-based evidence. These approaches reduce hallucination but depend heavily on retrieval quality and, in supervised variants, on labeled trajectories. Our setting differs fundamentally by treating a knowledge graph as the environment: actions are typed and executable against the KB schema, observations are structure-grounded entity sets rather than text passages, and step-wise executability can be validated programmatically rather than inferred from unstructured documents.
@@ III Preliminaries
KBQA Task. Given a natural language question $q$, the KB $\mathcal{K}$, and a set of topic entities $E_{q}\subseteq\mathcal{E}$ mentioned in $q$, the goal of Knowledge Base Question Answering (KBQA) is to produce an answer set $\mathcal{A}_{q}\subseteq\mathcal{E}$ that correctly responds to the question. Following prior work [luo2025kbqa], we assume that entity mentions in $q$ are already linked to the KB and the corresponding topic entities $E_{q}$ are given as input. In classic semantic-parsing based KBQA, this task is realized by generating a logical form (e.g., SPARQL or S-Expression) in one shot and executing it against the KB. In contrast, our framework rephrases the task as learning a multi-step interaction policy.
@@
Policy Optimization Objective. Our goal is to learn a policy that produces trajectories leading to correct answers. The action $a_{t}$ is grounded into an S-Expression fragment [ArcaneQA, GrailQA] and executed by the executor over $\mathcal{K}$. Let $R(\tau)$ denote the cumulative reward of a trajectory. From the reinforcement learning perspective, we optimize the expected return:
@@
where $r_{\text{outcome}}$ reflects answer correctness and $r_{\text{format}}$ rewards valid S-Expression structure. Unlike methods [luo2025kbqa] that require test-time search (e.g., MCTS), our approach trains a single policy end-to-end that directly generates high-quality trajectories at inference time without additional search overhead.
@@ IV-A Prompt and System Workflow
Our system is a multi-turn agent system inspired by the ReAct paradigm [ReAct]. At each turn, the LLM emits one or more actions to interact with the KB environment, and the environment returns the corresponding observations. After multiple turns of KB exploration, the model outputs the final answer.
@@ IV-A1 Prompting Template for Action-Based Reasoning
Table I shows the prompting template used to elicit action-based reasoning from the LLM. The template structures the model’s output into three parts in an iterative fashion: first, a reasoning process (<think>...</think>), similar to Chain-of-Thought prompting [CoT], then a Knowledge Graph exploration action (<action>...</action>, e.g., Find_relation, Merge), and finally the answer (<answer>...</answer>). Crucially, we only impose structural constraints on the output format, not on the reasoning content. This design choice ensures that the model learns to reason adaptively through RL, rather than mimicking template-driven patterns as in prior work [luo2025kbqa].
@@ IV-A2 Action Space
Prior semantic parsing approaches to KBQA [ArcaneQA, ChatKBQA, RnG-KBQA] typically require the model to emit a full, nested S-expression in a single pass. This design is notoriously brittle: a single token-level error (e.g., a typo in a relation name or a mismatched parenthesis) can render the entire program unexecutable and cause the query to fail.
Following the recent KBQA-o1 framework [luo2025kbqa], we instead adopt a compact, discrete action space that decomposes logical-form construction into a sequence of simple, verifiable steps. Concretely, each action corresponds to an atomic operation over the evolving logical expression, such as extending from an entity along a relation (Find_relation), intersecting two partial expressions (Merge), or applying aggregation and comparison operators (Order, Compare, Count, Time_constraint). As summarized in Table II, every action is defined by (i) its arguments, (ii) a target functional update on the current expression (e.g., JOIN, AND, ARG, CMP, TC, COUNT), and (iii) the corresponding S-expression template.
Operationally, the agent does not generate the complete program at once. Instead, starting from candidate entities detected in the question, it emits one or more actions at each turn, observes the execution results against the KB, and then decides the next action based on this feedback. This interleaved generation–execution process, inherited from related iterative retrieval methods [trivedi2022interleaving, jiang2023active], improves robustness in two ways: the environment can validate and correct individual actions (e.g., via schema-aware relation retrieval), and errors are localized to specific steps rather than invalidating the entire program.
Actions are converted into an S-Expression list, then translated into SPARQL queries [SPARQL] executed against the KB. The resulting observations are appended to the dialogue state visible to the model. This workflow mitigates the fragility of string-based S-expression program generation and lowers the error rate in actions produced by the LLM.
@@ IV-A3 Relation Retrieval and Confidence Gating
LLM-proposed relations can be noisy or ambiguous due to the well-known hallucination problem [hallucination]. To mitigate this, we introduce the Relation Retrieval and Confidence Gating (RRCG) module. The RRCG module acts as a validation layer, verifying the agent’s proposed textual relation before execution.
Let $r_{\text{agent}}$ be the original textual relation proposed by the agent for the current entity $e_{c}$. Let $R(e_{c})$ be the set of all neighboring schema relations of $e_{c}$ in the knowledge base. The core of the RRCG module is a similarity function $Sim(\cdot,\cdot)$, implemented using dense retrieval techniques [karpukhin2020dense, dense_retrieval], which scores $r_{\text{agent}}$ against every schema relation $r_{s}\in R(e_{c})$. We define $s_{\text{max}}=\max_{r_{s}\in R(e_{c})}Sim(r_{\text{agent}},r_{s})$ as the highest similarity score, with $r_{s}^{*}=\arg\max_{r_{s}\in R(e_{c})}Sim(r_{\text{agent}},r_{s})$ being the best-matching schema relation.
@@ IV-B Rejection Sampling and Supervised Fine-Tuning Warm-Start
The key insight behind RRS is that successful KBQA trajectories must align with executable action sequences. Standard rejection sampling [RFT] from raw prompts suffers from very low acceptance rates due to the task complexity and the base LLM’s weak zero-shot ability on structured KB queries. Simply increasing sampling temperature or budget yields diminishing returns, as most generated trajectories contain hallucinated relations or malformed S-Expressions.
RRS addresses this by providing the model with a reference action sequence—extracted from the gold S-Expression—as implicit guidance during generation. This approach is inspired by rationalization techniques in STaR [STaR], where hints are provided when the model fails, but we extend it to agentic settings with environmental interaction. This constraint forces the model to: ❶ Ground reasoning in execution: The model must justify why each reference action leads toward the correct answer, rather than fabricating post-hoc explanations. ❷ Learn action-observation correspondence: By observing the actual KB feedback for each ground-truth action, the model internalizes the mapping between actions and their environmental consequences.
@@ IV-B2 SFT Training
The SFT process fine-tunes the base LLM on $S_{\text{RRS}}$. Following best practices in instruction tuning [ouyang2022training], we compute the loss only on assistant-visible tokens (i.e., <think> reasoning and <action> blocks); tool messages (<information> segments) are masked from the loss and serve only as context. This selective masking ensures the model learns to generate actions and reasoning, not to memorize environmental feedback. The resulting SFT checkpoint initializes the policy for subsequent GRPO optimization.
@@ IV-C Reinforcement Learning Optimization
The policy is further refined via Reinforcement Learning [kaelbling1996reinforcement], optimizing a composite reward signal using our GRPO algorithm [shao2024deepseekmath].

### v2

@@ Abstract
Knowledge Base Question Answering (KBQA) challenges models to bridge the gap between natural language and strict knowledge graph schemas by generating executable logical forms. While Large Language Models (LLMs) have advanced this field, current approaches often struggle with a dichotomy of failure: they either generate hallucinated queries without verifying schema existence or exhibit rigid, template-based reasoning that mimics synthesized traces without true comprehension of the environment. To address these limitations, we present KBQA-R1, a framework that shifts the paradigm from text imitation to interaction optimization via Reinforcement Learning. Treating KBQA as a multi-turn decision process, our model learns to navigate the knowledge base using a list of actions, leveraging Group Relative Policy Optimization (GRPO) to refine its strategies based on concrete execution feedback rather than static supervision. Furthermore, we introduce Referenced Rejection Sampling (RRS), a data synthesis method that resolves cold-start challenges by strictly aligning reasoning traces with ground-truth action sequences. Extensive experiments on WebQSP, GrailQA, and GraphQuestions demonstrate that KBQA-R1 achieves state-of-the-art performance, effectively grounding LLM reasoning in verifiable execution. Code and data are available at https://github.com/sunxin000/KBQA-R1 . Project page: https://sunxin000.github.io/KBQA-R1/.
@@ I Introduction
Despite significant progress in applying LLMs to KBQA, existing methodologies can be categorized into three paradigms, each with distinct limitations. The first category comprises End-to-end Approaches (e.g., KB-BINDER [19], KB-Coder [25], ChatKBQA [21]), which generate entire logical forms in a single pass. While computationally efficient, these methods operate without intermediate KB interaction, making them unable to verify the existence of schema elements during generation. This often leads to schema hallucinations—syntactically valid but semantically incorrect queries that reference non-existent relations or entities. The second category, Prompting-based Step-by-Step Approaches [37, 23], leverages powerful commercial model APIs with few-shot in-context learning to decompose complex queries into sequential reasoning steps. While these methods benefit from the strong reasoning capabilities of large-scale models, they lack task-specific training, resulting in suboptimal performance on domain-specific schema navigation and complex multi-hop queries. The third category, Supervised Agent Approaches [22], mitigates the above issues by fine-tuning models on reasoning traces synthesized from templates or heuristics. However, this paradigm risks superficial reasoning: since the training data is inherently formulaic, the model’s “thoughts” often reduce to template-driven action announcements (e.g., “At this step, we should find the relation…”) rather than genuine analysis of environmental feedback. The model declares what action to take without explaining why—it neither interprets the KB observations nor justifies its choices based on the question semantics. Such rigid patterns, while achieving local consistency, fail to generalize when novel schema structures or unexpected query patterns arise.
To address these limitations, we present KBQA-R1, an action-centric reinforcement learning framework that shifts the paradigm from text imitation to interaction optimization. Instead of generating raw query code, KBQA-R1 operates within a well-defined action space, treating KBQA as a multi-turn Markov Decision Process (MDP). At each step, the model selects an action (e.g., Find_Relation, Merge), observes concrete feedback from the KB execution engine, and dynamically adjusts its reasoning trajectory. Crucially, because the policy is optimized via Group Relative Policy Optimization (GRPO) [32] with outcome-based rewards rather than imitation loss, the model is incentivized to develop adaptive reasoning—genuinely analyzing observations and justifying action choices—rather than memorizing fixed reasoning templates. This closed-loop interaction grounds decisions in verifiable outcomes and enables the model to discover effective reasoning strategies through environmental exploration.
@@ II Related Work
Knowledge Base Question Answering (KBQA). Before the rise of LLMs, KBQA studies are commonly categorized into information-retrieval-based (IR-based) methods [36, 35, 54, 8, 28] and semantic-parsing-based (SP-based) methods [47, 33, 6, 55]. With LLMs, three paradigms have emerged: (i) end-to-end approaches that directly generate logical forms via in-context learning or fine-tuning [19, 25, 21, 11]; (ii) step-by-step (agentic) approaches that interleave reasoning with graph exploration and tool use [4, 10, 37, 23, 43, 2, 12]; and (iii) search-augmented approaches that leverage tree search algorithms such as Monte Carlo Tree Search (MCTS) for systematic exploration [22].
@@
While MCTS-based methods like KBQA-o1 [22] achieve strong performance through heuristic exploration, they exhibit two key limitations. First, they incur significant computational overhead from multiple rollouts per query and require separate policy and reward models during inference. Second, their reasoning traces are often template-driven (e.g., “At this step, we should find the relation…”) rather than genuinely analytical—the model announces what action to take without explaining why based on observations. In contrast, we train a single policy via RL with outcome-based rewards, encouraging the model to develop adaptive reasoning that analyzes environmental feedback and justifies action choices, while eliminating test-time search overhead.
@@
LLMs, tool use, and agentic reasoning. Chain-of-Thought (CoT) prompting improves reasoning by eliciting intermediate steps [41]; ReAct [46] interleaves “think” and “act” to ground reasoning in environment feedback; and heuristic search has been applied to agent traces (e.g., MCTS-style selection in [7] and tree-structured deliberation in [45]). Recent graph-augmented approaches such as Plan-on-Graph [2] incorporate self-correcting mechanisms with dynamic memory for adaptive planning on knowledge graphs. While these methods expand the search space or stabilize multi-step reasoning, free-form thoughts can overfit prompt templates and do not guarantee executability. We keep the interleaved think-act design but require typed, schema-aware actions with validators and an executor, turning traces into verifiable computations rather than narrative justifications.
@@
Retrieval-augmented generation and search-as-a-tool. Classical RAG pipelines retrieve text snippets and feed them to the model for generation [18]. Recent work moves toward search-as-a-tool, prompting or training LLMs to issue search calls and iterate [40, 46, 29, 20, 14]. GraphRAG approaches [9, 24] further integrate graph retrieval with LLM reasoning, enabling tighter coupling between structured knowledge and text-based evidence. These approaches reduce hallucination but depend heavily on retrieval quality and, in supervised variants, on labeled trajectories. Our setting differs fundamentally by treating a knowledge graph as the environment: actions are typed and executable against the KB schema, observations are structure-grounded entity sets rather than text passages, and step-wise executability can be validated programmatically rather than inferred from unstructured documents.
@@ III Preliminaries
KBQA Task. Given a natural language question $q$, the KB $\mathcal{K}$, and a set of topic entities $E_{q}\subseteq\mathcal{E}$ mentioned in $q$, the goal of Knowledge Base Question Answering (KBQA) is to produce an answer set $\mathcal{A}_{q}\subseteq\mathcal{E}$ that correctly responds to the question. Following prior work [22], we assume that entity mentions in $q$ are already linked to the KB and the corresponding topic entities $E_{q}$ are given as input. In classic semantic-parsing based KBQA, this task is realized by generating a logical form (e.g., SPARQL or S-Expression) in one shot and executing it against the KB. In contrast, our framework rephrases the task as learning a multi-step interaction policy.
@@
Policy Optimization Objective. Our goal is to learn a policy that produces trajectories leading to correct answers. The action $a_{t}$ is grounded into an S-Expression fragment [6, 5] and executed by the executor over $\mathcal{K}$. Let $R(\tau)$ denote the cumulative reward of a trajectory. From the reinforcement learning perspective, we optimize the expected return:
@@
where $r_{\text{outcome}}$ reflects answer correctness and $r_{\text{format}}$ rewards valid S-Expression structure. Unlike methods [22] that require test-time search (e.g., MCTS), our approach trains a single policy end-to-end that directly generates high-quality trajectories at inference time without additional search overhead.
@@ IV-A Prompt and System Workflow
Our system is a multi-turn agent system inspired by the ReAct paradigm [46]. At each turn, the LLM emits one or more actions to interact with the KB environment, and the environment returns the corresponding observations. After multiple turns of KB exploration, the model outputs the final answer.
@@ IV-A1 Prompting Template for Action-Based Reasoning
Table I shows the prompting template used to elicit action-based reasoning from the LLM. The template structures the model’s output into three parts in an iterative fashion: first, a reasoning process (<think>...</think>), similar to Chain-of-Thought prompting [41], then a Knowledge Graph exploration action (<action>...</action>, e.g., Find_relation, Merge), and finally the answer (<answer>...</answer>). Crucially, we only impose structural constraints on the output format, not on the reasoning content. This design choice ensures that the model learns to reason adaptively through RL, rather than mimicking template-driven patterns as in prior work [22].
@@ IV-A2 Action Space
Prior semantic parsing approaches to KBQA [6, 21, 47] typically require the model to emit a full, nested S-expression in a single pass. This design is notoriously brittle: a single token-level error (e.g., a typo in a relation name or a mismatched parenthesis) can render the entire program unexecutable and cause the query to fail.
Following the recent KBQA-o1 framework [22], we instead adopt a compact, discrete action space that decomposes logical-form construction into a sequence of simple, verifiable steps. Concretely, each action corresponds to an atomic operation over the evolving logical expression, such as extending from an entity along a relation (Find_relation), intersecting two partial expressions (Merge), or applying aggregation and comparison operators (Order, Compare, Count, Time_constraint). As summarized in Table II, every action is defined by (i) its arguments, (ii) a target functional update on the current expression (e.g., JOIN, AND, ARG, CMP, TC, COUNT), and (iii) the corresponding S-expression template.
Operationally, the agent does not generate the complete program at once. Instead, starting from candidate entities detected in the question, it emits one or more actions at each turn, observes the execution results against the KB, and then decides the next action based on this feedback. This interleaved generation–execution process, inherited from related iterative retrieval methods [40, 13], improves robustness in two ways: the environment can validate and correct individual actions (e.g., via schema-aware relation retrieval), and errors are localized to specific steps rather than invalidating the entire program.
Actions are converted into an S-Expression list, then translated into SPARQL queries [27] executed against the KB. The resulting observations are appended to the dialogue state visible to the model. This workflow mitigates the fragility of string-based S-expression program generation and lowers the error rate in actions produced by the LLM.
@@ IV-A3 Relation Retrieval and Confidence Gating
LLM-proposed relations can be noisy or ambiguous due to the well-known hallucination problem [56]. To mitigate this, we introduce the Relation Retrieval and Confidence Gating (RRCG) module. The RRCG module acts as a validation layer, verifying the agent’s proposed textual relation before execution.
Let $r_{\text{agent}}$ be the original textual relation proposed by the agent for the current entity $e_{c}$. Let $R(e_{c})$ be the set of all neighboring schema relations of $e_{c}$ in the knowledge base. The core of the RRCG module is a similarity function $Sim(\cdot,\cdot)$, implemented using dense retrieval techniques [16, 57], which scores $r_{\text{agent}}$ against every schema relation $r_{s}\in R(e_{c})$. We define $s_{\text{max}}=\max_{r_{s}\in R(e_{c})}Sim(r_{\text{agent}},r_{s})$ as the highest similarity score, with $r_{s}^{*}=\arg\max_{r_{s}\in R(e_{c})}Sim(r_{\text{agent}},r_{s})$ being the best-matching schema relation.
@@ IV-B Rejection Sampling and Supervised Fine-Tuning Warm-Start
The key insight behind RRS is that successful KBQA trajectories must align with executable action sequences. Standard rejection sampling [52] from raw prompts suffers from very low acceptance rates due to the task complexity and the base LLM’s weak zero-shot ability on structured KB queries. Simply increasing sampling temperature or budget yields diminishing returns, as most generated trajectories contain hallucinated relations or malformed S-Expressions.
RRS addresses this by providing the model with a reference action sequence—extracted from the gold S-Expression—as implicit guidance during generation. This approach is inspired by rationalization techniques in STaR [53], where hints are provided when the model fails, but we extend it to agentic settings with environmental interaction. This constraint forces the model to: ❶ Ground reasoning in execution: The model must justify why each reference action leads toward the correct answer, rather than fabricating post-hoc explanations. ❷ Learn action-observation correspondence: By observing the actual KB feedback for each ground-truth action, the model internalizes the mapping between actions and their environmental consequences.
@@ IV-B2 SFT Training
The SFT process fine-tunes the base LLM on $S_{\text{RRS}}$. Following best practices in instruction tuning [26], we compute the loss only on assistant-visible tokens (i.e., <think> reasoning and <action> blocks); tool messages (<information> segments) are masked from the loss and serve only as context. This selective masking ensures the model learns to generate actions and reasoning, not to memorize environmental feedback. The resulting SFT checkpoint initializes the policy for subsequent GRPO optimization.
@@ IV-C Reinforcement Learning Optimization
The policy is further refined via Reinforcement Learning [15], optimizing a composite reward signal using our GRPO algorithm [32].